OpenStudy (anonymous):

How do you factor completely?: -t^2-9t+1

ganeshie8 (ganeshie8):

you cannot factor it in integers, i think we call such polynomials as prime polynomials

OpenStudy (fibonaccichick666):

so, let us use real numbers :)

OpenStudy (anonymous):

PRIME was not a correct response

OpenStudy (fibonaccichick666):

what have you tried @v_nessa_87 ?

OpenStudy (anonymous):

Factoring out -1, trying to apply perfect square trinomials method but the -9t is throwing me off.

OpenStudy (fibonaccichick666):

so, let's see: if I said this is of the form \(ax^2+bx+c\) what immediately jumps to mind?

OpenStudy (anonymous):

FOIL

OpenStudy (fibonaccichick666):

what else?

OpenStudy (anonymous):

That's it. I tried rewriting it like this: -1(t2+(3^2)t-(1^2)) but still don't think that's the right way to go.

OpenStudy (fibonaccichick666):

just wait, you're ignoring the obvious

OpenStudy (fibonaccichick666):

if you had \(ax^2+bx+c=0\) how would you get the x values?

OpenStudy (anonymous):

Nothing multiplies to -1 and adds to +9.

OpenStudy (campbell_st):

well I think you need to use the discriminant to investigate if it can be factored the discriminant is \[\Delta = b^2 - 4ac\]

OpenStudy (fibonaccichick666):

That doesn't matter

OpenStudy (fibonaccichick666):

There is a formula, you probably have memorized

OpenStudy (fibonaccichick666):

usually starts with a Q

OpenStudy (anonymous):

I don't want the value of T, I want to factor.

OpenStudy (anonymous):

Isn't the q form for values?

OpenStudy (fibonaccichick666):

Factoring and finding the value go hand in hand

OpenStudy (campbell_st):

if \[\Delta > 0\] and is a square number then the quadratic has 2 unequal roots, and it can be factored

OpenStudy (fibonaccichick666):

and I do not know a Q form for values. No, like seriously, in 6/8th gradish how would you have found the values of x in \(ax^2+bx+c=0\) What formula would you have used?

OpenStudy (campbell_st):

so use the discriminant so you don't waste you time trying to factor

OpenStudy (anonymous):

quadratic formula. i abbreviated: q.uadratic form.ula

OpenStudy (fibonaccichick666):

YEssss

OpenStudy (fibonaccichick666):

so, apply it here, once you do

OpenStudy (fibonaccichick666):

you get two values, those are your second terms in your factoring

OpenStudy (fibonaccichick666):

I recommend pulling out the -1 first though

OpenStudy (fibonaccichick666):

you will get a real answer

OpenStudy (campbell_st):

so given you are asked to factor... the quadratic is prime since \[\Delta = (-9)^2 - 4 \times (-1) \times 1 = 85\] not a square number

OpenStudy (fibonaccichick666):

You do not need an integer solution to factor... I don't know why you are implying that

OpenStudy (anonymous):

-9/2 & sqrt(85)/2

OpenStudy (anonymous):

-(t-9/2)(t+sqrt(85/2) ?

OpenStudy (fibonaccichick666):

actually no, I don't buy it

OpenStudy (fibonaccichick666):

you shouldn't have a 9 for one and a \(\sqrt{85}\) for the other

OpenStudy (fibonaccichick666):

check your quadratic answers, but you set it up well

OpenStudy (anonymous):

The only reason I am hesitant using the quadratic formula is generally assignments don't have you use skills not yet taught in previous lessons. The quadratic formula isn't introduced for another two lessons ahead of this assignment.

OpenStudy (fibonaccichick666):

well, I will say it is how I'd do this. Otherwise they want you to complete the square

OpenStudy (fibonaccichick666):

(I don't like completing the square if I don't have to

ganeshie8 (ganeshie8):

why do you think PRIME is not a correct response ? @v_nessa_87

OpenStudy (anonymous):

because I submitted that as the answer and it returned incorrect

ganeshie8 (ganeshie8):

try submitting this \[-t^2-9t+1 = -1(t^2+9t-1)\] there is nothing much you can do here really..

ganeshie8 (ganeshie8):

(ofcourse in integers)

OpenStudy (anonymous):

Oh man, I can't believe all it was was factoring out -1. If it had been t^2+9t-1, would it have been prime? That drives me nuts.

ganeshie8 (ganeshie8):

lol does that work ?

OpenStudy (anonymous):

Yeah, it worked... *sigh*

ganeshie8 (ganeshie8):

yaaay! to me, t^2+9t-1 is as much prime as -t^2-9t+1 is prime factoring out -1 is silly but yeah just give the grader whatever it wants haha!

OpenStudy (fibonaccichick666):

oh jeeze, these online assignments are so dumb

OpenStudy (anonymous):

^agreed

OpenStudy (anonymous):

thanks all for your help!